(5 marks) Algorithm performance Explain a) (4 marks) Give the order of magnitude e Ofor the following algorithm. why your answer is correct. GET VALUES for A1, A2, ...3 An, and B1, B Bn Get value of n set i equal to 1 i 1 for each of the n values in A DO WHILE (is n) set j equal to 1 i 1 DO WHILE G S n) Do n times IF Ai Bj THEN PRINT Found it! increment j j j+1 END DO WHILE increment i END DO WHILE (1 mark) Explain what the algorithm does.
Answer #1

For each value of A, it is searching in B. So basically this algorithm finding the comman elements.

Here we have two do while loop. One is nested in other.

Since size of A and B is n.

So, T(n) = 1 + 2+ 3+ ...+n = n(n+1)/2 = O(n^2)